Sofia files married filing joint and has the following amounts in 2018:

Net income reported on Sch. C (net of $120,000 wages, no significant assets)

$500,000

Net capital gains

$30,000

Taxable income before Qualifying Business Income Deduction

$600,000

What is the amount of the Qualified Business Income Deduction?

A. $114,000
B. None of the other choices are correct.
C. $120,000
D. $0.00
E. $60,000
F. $100,000
Answer #1

Option E

$60000

QBI deduction starts to phase out in case of taxable income exceeds $315000 for married filing jointly. In that case QBID is lesser of

20% of qualified business income

OR

Greater of 50% of the W-2 wages related to the qualified trade or business or sum of 25% of the W-2 wages related to the qualified trade or business, + 2.5% of the UBIA of qualified property

Greater of 50% of the W-2 wages related to the qualified trade or business or sum of 25% of the W-2 wages related to the qualified trade or business, + 2.5% of the UBIA of qualified property = (120000*50%) or (120000*25%)+(2.5%*0) = 60000 or 30000

Thus, it is $60000

Now lesser of 20% of qualified business income (500000*20%) = $100000

Or

$60000

Thus, the amount of the Qualified Business Income Deduction is $60000